OpenStudy (anonymous):

One spinner has 24 equal sections and 8 are red. A second spinner has 32 equal sections and 8 are red. What is the probability that red will be spun on both spinners?

OpenStudy (kropot72):

The outcomes of the two spins are independent. Therefore the intersection of theprobability of red being spun on both is the product of the individual probabilities of red. P(red on both) = (8/24) * (8/32) = (1/3) * (1/4)

OpenStudy (anonymous):

thank you!

OpenStudy (kropot72):

You're welcome :)